In how many ways can the letters of the word PATLIPUTRA be arranged, so that the relative order of vowels and consonants do not alter?

Grade:12th pass

1 Answers

Arun
6 years ago
The word PATALIPUTRA has total 11 characters. Out of them, 
5 are Vowels (A A I U A) occupying positions 2,4,6,8,11
6 are consonants (P T L P T R) occupying positions 1,3,4,5,8,9
 
Relative order of vowels and consonants cannot be changed.
 i.e., position of vowels and consonants cannot be changed

5 vowels can be arranged in 5! / 3! ways  (we divide by 3! as there are 3 'A')

6 consonants can be arranged in 6! / 2!×2! ways (we divide by (2! × 2!) as there are 2 'P' and 2 'T')

Total number of ways = (5!/ 3!)×(6!/2!×2!)=20×180=3600